OpenStudy (alexwee123):

given (x-2)^2/16 + y^2/12 =1, an ellipse, how would you find the polar equation of this?

OpenStudy (anonymous):

by sub x-h=a*cos(tetha),y=b*sin(tetha)

OpenStudy (alexwee123):

can you show me your steps?

OpenStudy (anonymous):

by comparing h=2,a=4,b=12^(1/2)and polar eqs are x-2=4*cos(tetha),y=12^(1/2)sin(tetha)

OpenStudy (alexwee123):

hmm can you put it into standard orientation of this conic?